undeadknight26 (undeadknight26):

Jason had $32. He spent all the money on buying 4 CDs for $x each and 2 magazines for $y each. If Jason would have bought 5 CDs and 2 magazines, he would have run short of $4.

undeadknight26 (undeadknight26):

The student's conclusion is incorrect because the solution to the system of equations 4x + 2y = 4 and 5x + 2y = 36 is (9, 4). The student's conclusion is incorrect because the solution to the system of equations 4x + 2y = 32 and 5x + 2y = 36 is (4, 8). The student's conclusion is correct because the solution to the system of equations 4x + 2y = 4 and 5x + 2y = 36 is (5, 6). The student's conclusion is correct because the solution to the system of equations 4x + 2y = 32 and 5x + 2y = 36 is (5, 6).

undeadknight26 (undeadknight26):

@PixieDust1 @amistre64 @alyssajobug @charlotte123 @bahrom7893 @Hannah28 @LilySwan @kawaiicat123

OpenStudy (anonymous):

why dont you graph it

OpenStudy (pixiedust1):

This question has been closed... are you sure your not still having trouble with it? I do not see any people walking you through it.

undeadknight26 (undeadknight26):

im good thx.